OpenStudy (anonymous):

The express train traveled at a rate of speed that was three times as fast as the freight train. The express train traveled 630 miles in just one hour less than it took the freight train to travel 360 miles. Find the rate of speed and the time for both the express train and the freight train.

OpenStudy (wasiqss):

we can make two equations here

OpenStudy (wasiqss):

let speed of freight train be X km/h let speed of express train be 3X km/h

OpenStudy (wasiqss):

you there

OpenStudy (anonymous):

Yes, just trying to work it out.

OpenStudy (wasiqss):

did you get the first step

OpenStudy (anonymous):

That's really still very confusing. Could you write the formula?

OpenStudy (wasiqss):

i mean speed for freight train = X miles/hour speed for express traing= 3X miles/hour

OpenStudy (anonymous):

Yes, I get that. But I mean how would I be able to work that out properly?

OpenStudy (wasiqss):

ok i would tell you

OpenStudy (wasiqss):

suppose the frieght train took Y hours to travel 360 miles ok?

OpenStudy (anonymous):

Right

OpenStudy (wasiqss):

then Y= distance/speed alright?

OpenStudy (wasiqss):

where y =time

OpenStudy (wasiqss):

distance=360, speed = X ok? hence Y=360/X

OpenStudy (anonymous):

So the speed of the express is 150 and the speed of the freight is 450. But I can't figure out the times.

OpenStudy (wasiqss):

dont be quick to get answers try to go step by step

OpenStudy (wasiqss):

:D

OpenStudy (anonymous):

Well I did ReTe = 360 | RfTf = 630 | Rf = 3Re | Tf = Te - 1 Then I substituted and that's how I found my answer for the speeds. But then when I went to work on the times, I can't seem to get it to fit together.

OpenStudy (wasiqss):

your answers are right? and method is right

OpenStudy (anonymous):

Yes, those first two are correct.

OpenStudy (wasiqss):

so now what you are stuck in

OpenStudy (anonymous):

Well following the answers I got... The times would have to be 1.4 and 2.4 - Right?

OpenStudy (wasiqss):

lol you made a dumb mistake :P the speed for frieght train=150, and for express it is 450 , now you ll get answer

OpenStudy (anonymous):

Eh... 1.25 and 4.2?

OpenStudy (wasiqss):

630/450= 1.4 right, that is for express now 360/150=2.4 right, that is for freight, now you can see diffrence is of an hour

OpenStudy (wasiqss):

you got it

OpenStudy (anonymous):

Ha! Sorry, yes, I got it now. Thank-you.

OpenStudy (wasiqss):

lol hard earned medal it was :D